### v1

# TabXEval: Why this is a Bad Table? An eXhaustive Rubric for Table Evaluation

Abstract
Evaluating tables qualitatively & quantitatively presents a significant challenge, as traditional metrics often fail to capture nuanced structural and content discrepancies. To address this, we introduce a novel, methodical rubric integrating multi-level structural descriptors with fine-grained contextual quantification, thereby establishing a robust foundation for comprehensive table comparison. Building on this foundation, we propose TabXEval, an eXhaustive and eXplainable two-phase evaluation framework. TabXEval initially aligns reference tables structurally via TabAlign & subsequently conducts a systematic semantic and syntactic comparison using TabCompare; this approach clarifies the evaluation process and pinpoints subtle discrepancies overlooked by conventional methods. The efficacy of this framework is assessed using TabXBench, a novel, diverse, multi-domain benchmark we developed, featuring realistic table perturbations and human-annotated assessments. Finally, a systematic analysis of existing evaluation methods through sensitivity-specificity trade-offs demonstrates the qualitative and quantitative effectiveness of TabXEval across diverse table-related tasks and domains, paving the way for future innovations in explainable table evaluation.
TabXEval: Why this is a Bad Table?
An eXhaustive Rubric for Table Evaluation
thanks: These authors contributed equally to this work.Vihang Pancholi [^1] footnotemark: 1Jainit Bafna211footnotemark: 1Tejas Anvekar1 Manish Shrivastava2thanks: Primary supervisor of this work.Vivek Gupta1 1Arizona State University  2IIIT Hyderabad {vpancho1,tanvekar,vgupta140}@asu.edu {jainit.bafna,m.shrivastava}@research.iiit.ac.in
1 Introduction
Tables are a ubiquitous data format across a multitude of domains from financial reporting and medical data analysis to scientific research and knowledge bases. Each of these fields confronts its own unique challenges: financial tables must maintain rigorous numerical precision and consistency, while medical tables demand exactitude in categorical and temporal data; both scenarios underscore the difficulty of comparing tables when discrepancies may arise from subtle semantic, structural, or formatting variations. Traditional evaluation metrics originally designed for text generation, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), METEOR Banerjee and Lavie (2005), and chrF Popović (2015), rely on surface-level n–gram overlaps and fail to capture these multifaceted discrepancies. Even embedding-based approaches like BERTScore Zhang et al. (2020), which excel in measuring semantic similarity in text-to-text tasks, fall short when applied to the intricacies of tabular data.
Recent efforts have sought to address these limitations. Benchmarks such as StructBenchGu et al. (2024) and TanQAkhtar et al. (2025) introduce complex structural challenges and query-oriented tasks that push the boundaries of table generation. In parallel, methods like the “Is this a bad table?” metric Ramu and Bandyopadhyay (2024) decompose tables into atomic natural language statements to assess errors in a more granular fashion. However, these approaches often compromise either on the explainability or the granularity needed for robust table evaluation. Moreover, in contrast to domains like image captioning where datasets and benchmarks such as THumB Kasai et al. (2022) provide comprehensive human evaluations to validate metrics table evaluation lacks a comparable level of benchmark diversity and human-aligned assessment.
To overcome these challenges, we propose a novel evaluation framework, TabXEval, which is built upon a comprehensive, multi-level rubric that integrates structural descriptors with fine-grained contextual quantification. TabXEval operates in two distinct phases: TabAlign first establishes a precise baseline alignment of table elements via both deterministic and LLM-based methods, and TabCompare subsequently conducts a detailed semantic and syntactic analysis of aligned cells. This two-phase approach enables us to capture both coarse (table-level) and fine-grained (cell-level) discrepancies that conventional metrics overlook.
Complementing our evaluation method is TabXBench, a synthetic yet diverse benchmark designed to mimic real-world table perturbations across multiple domains. By incorporating human-annotated assessments aligned with our proposed rubric, TabXBench serves as a crucial testbed for quantifying sensitivity, specificity, and overall alignment with human judgment addressing a significant gap in current evaluation paradigms.
  • We introduce the first rubric integrating multi-level structural descriptors and fine-grained contextual quantification for robust table comparisons.
  • We propose TabXEval, a two-phase LLM-based table evaluation method that aligns reference tables structurally and compares them semantically and syntactically via our rubric.
  • We construct TabXBench, a diverse benchmark derived from multi-domain datasets, validating evaluation metrics through structured perturbations and human assessments.
  • We analyze the strengths and weaknesses of existing evaluation methods via Sensitivity-Specificity Trade-off.
  • We demonstrate TabXEval’s qualitative and quantitative effectiveness in text-to-table generation, fostering explainable automatic evaluation.
Our TabXBench dataset and TabXEval source code are available at https://coral-lab-asu.github.io/tabxeval/.
2 Related Works
Early work in text-to-table generation predominantly leveraged datasets originally designed for table-to-text tasks such as RotoWireWiseman et al. (2017), E2ENovikova et al. (2017), WikiBioLebret et al. (2016), and WikiTableTextPasupat and Liang (2015) which, while pioneering, suffer from inherent limitations. These datasets are often characterized by a lack of structural diversity and a tendency to produce hallucinated responses, thereby compromising both the factual accuracy and robustness of generated tables. Recent benchmarks, including StrucBenchGu et al. (2024)and TanQAkhtar et al. (2025), have emerged to address these shortcomings by introducing more complex structural challenges and query-oriented generation tasks that emphasize multi-hop reasoning and multi-source data integration. However, these new benchmarks remain under-explored, underscoring the need for further work to validate their effectiveness and bridge the gap between theoretical advancements and practical performance.
Traditional evaluation metrics for generated text such as BLEU Papineni et al. (2002), ROUGELin (2004), METEORBanerjee and Lavie (2005), and chrFPopović (2015) primarily assess surface-level n–gram overlaps between candidate and reference texts. To capture deeper semantic similarity, embedding-based methods like BERTScoreZhang et al. (2020) were developed; yet, when applied to table evaluation, they fall short. In the context of tables, key structural elements such as row and column alignment, header consistency, and hierarchical relationships are critical. Metrics that operate on a token or cell–level basis, including Exact Match, PARENTDhingra et al. (2019), and even BERTScoreZhang et al. (2020), fail to account for the interdependent and contextual nature of table cells, leading to an incomplete assessment of both semantic and structural fidelity.
Recent work in StructBench Gu et al. (2024)introduces the H-score and p-score for table evaluation. The H-score measures hierarchical alignment and contextual consistency, while the p-score incorporates probabilistic confidence intervals to assess prediction reliability. These metrics address the limitations of token-based methods by focusing on structural coherence and robustness.
To overcome these limitations, recent research has explored hybrid approaches that combine structural decomposition with semantic evaluation. For example, the “Is this a bad table?” metric Ramu and Bandyopadhyay (2024) decomposes tables into natural language atomic statements and applies entailment-based scoring to capture both syntactic and semantic discrepancies in a holistic manner. Together, these efforts highlight an ongoing shift towards more exhaustive and explainable evaluation frameworks such as our proposed TabXEval that jointly consider both the structural alignment and contextual integrity of tables, thereby paving the way for more robust and generalizable table-to-table comparisons.
3 TabXEval
We establish a transparent evaluation protocol for reference table evaluation. Our rubrics and rules are developed through rigorous research survey and designed to increase the reliability of evaluation.
Figure 1: End-to-end schematic of TabXEval. (1) TabAlign aligns rows, columns, and cells using deterministic rules plus an LLM refinement loop. (2) TabCompare classifies each aligned cell as extra, missing, or partial and combines the counts with rubric weights $(\alpha,\beta,\gamma)$. This workflow populated the rubrics and outputs a table-level score and cell-level error trace, enabling fine-grained analysis.
3.1 Evaluation Rubric
To ensure consistent and fair structural evaluation, we introduce a comprehensive rubric that promotes clarity and transparency in reference-based table assessment. It quantifies correctness and information coverage through four key evaluation categories:
Structure Descriptor provides a coarse assessment at the table level capturing missing, extra, and matched information to reflect overall integrity.
Column Descriptor detects column data types and highlights missing or extra content. This enables fine-grained evaluation, especially for heterogeneous columns (e.g., dates, numbers, strings), avoiding uniform penalization.
Cell-Level Descriptor evaluates both semantic and syntactic correctness of cell values. Unlike prior methods Zhang et al. (2020); Ramu and Bandyopadhyay (2024); Gu et al. (2024), it correctly handles variant representations of the same data, reducing false penalties.
Granular Cell-Level Difference measures the degree of mismatch at the cell level, accounting for format shifts (e.g., meters to cm, years to full dates), and captures absolute differences crucial for fine-grained scoring.
3.2 Method
We propose TabXEval, a two-phase framework that combines deterministic rules and LLM-based analysis for robust and interpretable table evaluation. This design strikes a balance between precision (capturing exact matches) and flexibility (handling semantic or structural variations).
Phase 1: TabAlign
matches columns, rows, and cells between the reference and candidate tables. We begin with exact string matching to establish a precise baseline alignment. Next, we refine this alignment using an LLM to account for abbreviations, synonyms, and structural transformations (e.g., merged columns, row/column transpositions). Purely exact matching can be overly strict, missing semantically equivalent but syntactically different cells. The LLM-driven refinement ensures a more comprehensive alignment while preserving high precision. Finally, we get an output table which has a combination of strict and relaxed mapping as shown in  Figure 1.
Phase 2: TabCompare
performs a fine-grained evaluation of the aligned tables. From the refined alignment, we extract table-level statistics (e.g., missing/extra rows or columns) and focus on partially matched cells. These cells are compared in detail using LLM-generated “comparison tuples” as shown in Figure 1 which capture numeric, string, date/time, and unit mismatches. We also compute magnitudes of differences (e.g., converting months to days) for precise reporting of discrepancies. Table-level summaries alone cannot uncover subtle cell-level errors, such as unit mismatches or minor numeric discrepancies. By combining table-level statistics with granular cell comparisons, TabXEval yields a more reliable and transparent assessment of content fidelity.
Score
Our scoring function for TabXEval is defined as follows:
TabXEval $=\sum_{I\in\{\text{Missing},\,\text{Extra},\,\text{Partial}\}}\beta_{I}$
×$\quad\times\Bigl{(}\sum_{E\in\{\text{row},\,\text{column},\,\text{cell}\}}\alpha_{E}\,\tfrac{f_{E}}{N_{E}}\Bigr{)}\gamma_{p}\,.$
where $\beta_{I}$ is the weight assigned to each type of information error ($I$) such as Missing, Extra, and Partial; $\alpha_{E}$ is the weight for each entity type ($E$) including rows, columns, and cells; $f_{E}$ represents the number of correctly matched entities; and $N_{E}$ is the total number of entities in the ground truth.
For partial matches at the cell level, the modifier $\gamma_{p}$ is defined as:
$\gamma_{p}=\begin{cases}1,&\text{if no partial cell},\\[2.84526pt] \omega_{p}\,\bigl{\lvert}\tfrac{GT-Ref}{Ref}\bigr{\rvert},&\text{if partial cell detected}.\end{cases}$
This formulation captures the multi-level nature of table evaluation insipred by proposed rubric. First aggregating errors across different information types (Missing, Extra, Partial) via the outer summation, and then evaluating the correctness at various entity levels (row, column, cell) using the inner summation. The term $\gamma_{p}$ further refines the score by quantifying discrepancies in partially matched cells through a normalized absolute difference between the ground truth ($GT$) and the reference ($Ref$), ensuring that both coarse structural errors and fine-grained content differences are robustly accounted for in a single, interpretable metric.
Overall, TabXEval’s two-phase structure ensures that both coarse (table-level) and fine (cell-level) differences are captured, providing an adaptable and explainable approach to table evaluation.
3.3 TabXBench
Figure 2: Perturbation spectrum in TabXBench. The outer ring enumerates the frequency (numeric labels) of the 16 fine-grained perturbation types applied to reference tables. The inner ring groups these edits into three difficulty bands Easy (light green, $\approx$44%), Medium (blue, $\approx$34%), and Hard (red, $\approx$35%).
Evaluating table metrics across diverse domains and error types remains a significant challenge due to the limited scope of existing datasets, which often focus on a single domain or contain only select data types. To bridge this gap, we introduce TabXBench, a controlled multi-domain test bed that comprehensively captures real-world nuances of tabular data generation and evaluation.
Dataset # of Tables # Perturbation/Table Average
Rows Cols Cells
FINQA 13 5 13.08 5.54 71.00
TANQ 8 5 8.38 5.38 44.25
ROTOWIRE 12 5 12.08 14.08 165.67
FETAQA 7 5 15.86 5.14 88.14
WIKITABLES 5 5 16.40 3.60 53.60
WIKISQL 6 5 8.33 5.17 42.17
Table 1: Composition of the TabXBench corpus across six source datasets.
TabXBench is designed to rigorously assess the sensitivity and specificity of reference-based table evaluation methods. Unlike prior datasets that narrowly target specific tasks (e.g., finance or sports), our benchmark spans multiple domains (finance, sports, knowledge bases, and more) while systematically incorporating an extensive range of potential table perturbations. This diversity ensures that TabXBench captures common pitfalls such as missing rows/columns, reordered headers, unit mismatches, numeric discrepancies, and complex structural variations (e.g., row/column transposition).
TabXBench consists of 50 handpicked “clean” (reference) tables from six popular text-to-table and table QA datasets: RotoWireWiseman et al. (2017), TANQAkhtar et al. (2025), FetaQANan et al. (2021), FinQA Chen et al. (2021), WikiTable Pasupat and Liang (2015), and WikiSQLZhong et al. (2017) and the statistics can be found in Table 1. Each table was augmented with five distinct perturbations spanning over 16 error types, carefully curated to reflect common generation mistakes. The perturbations were first drafted with LLM assistance (e.g., reformatting numeric values, altering units, swapping headers) and subsequently validated by human experts for correctness and variety. We categorize the resulting perturbed tables into three difficulty levels (Easy, Medium, Hard) as demonstrated in Figure 2, ensuring coverage of both straightforward errors (e.g., minor typos) and complex structural manipulations (e.g., merged cells or shifted rows).
By offering a controlled yet diverse environment, TabXBench enables: Fine-Grained Analysis: Researchers can systematically evaluate how well metrics detect specific error types (e.g., unit mismatches vs. missing rows). Sensitivity-Specificity Trade-offs: The benchmark’s difficulty levels and variety of perturbations allow detailed insights into a method’s robustness and tolerance for minor vs. major errors. Realistic Scoring Correlations: It includes human annotations (aligned with our rubric in Section 3.1), enabling correlation studies that compare machine-generated scores to human judgments.
Finally, to illustrate TabXEval pragmatic value, we apply it to evaluate table outputs from three LLMs across four standard tasks - RotoWire, TANQ, WikiBio, and WikiTable highlighting how each evaluation method fares in a realistic setting. As a pioneering multi-domain resource, TabXBench thus provides a solid foundation for advancing research in reliable, explainable table evaluation.
4 Experiments
To validate efficacy of TabXEval, we conduct experiments using our synthetic dataset TabXBench. We report GPT-4oOpenAI et al. (2024) and LLaMA-3.3-INSTRUCT results for our framework TabXEval for both components TabAlign and TabCompare.
Baselines. Our evaluation compares TabXEval against a broad range of baselines, which we classify into deterministic and non-deterministic approaches. Deterministic metrics (e.g., Exact Match (EM), chrf, ROUGE-L) yield fixed outputs based on string- or character-level comparisons, ensuring reproducibility. In contrast, non-deterministic metrics (e.g., BERTScore, BLUERT, H-Score) leverage contextualized neural representations to capture subtle linguistic nuances albeit with potential variability. We also include two recent methods: P-Score, an LLM-based metric outputting scores on a 0–10 scale, and TabEval, an embedding-based method that unrolls tables using an LLM and computes entailment via RoBERTa-MNLI. For a fair comparison, we further propose a Direct-LLM baseline that is prompted with the same evaluation rubric detailed in Section 3.1.
LLMs. Throughout our experiments, we used GPT-4o, Gemini 2.0-flash, and LLaMA-3.3-Instruct 70B. All models were executed with identical sampling settings (default temperature, top-$k$, and top-$p$) unless specified otherwise. All our prompts for TabAlign, TabCompare and Direct-LLM basline are given in Figure 6, Figure 7, Figure 8.
4.1 Human correlation
For each sample in the ground truth set of TabXBench, two human evaluators evaluated the tables using our proposed rubrics and guidelines. For both Human and TabXEval, we present ground-truth and a randomly selected perturbation (out of 5) to fill the rubrics. Finally these rubrics are compared using both Pearson Sedgwick (2012) and Kendall’s Tau Sen (1968) correlation coefficients to quantify the degree of alignment between our method and human ratings. We observe very high correlation, i.e, $\mathbf{99.7\%}$ and $\mathbf{95.1\%}$ Pearson’s $\rho$ correlation for the Rubric Structure Descriptor and Cell Level Descriptor respectively. In contrast, Direct LLM based baseline correlates, only $\mathbf{30.6\%}$ and $\mathbf{40.6\%}$ Pearson’s $\rho$ correlation for the Rubric Structure Descriptor and Cell Level Descriptor respectively. Revealing that it fails to understand the rubric, and quantify the structural and contextual challenges in table evaluation, and hence is unable to correctly align with humans correlation. Similarly we report $\mathbf{99.1\%}$ and $\mathbf{92.8\%}$ human correlation using Kendall’s $\tau$. While the baseline is $\mathbf{30.7\%}$ and $\mathbf{55\%}$ on Structure Descriptor and Cell Level Descriptor rubric respectively. These results demonstrate that disentangling the alignment and comparison phases is critical for robust table evaluation as Direct method falls short even hen presented with evaluation rubrics. TabXEval’s correlation with human are consistent in capturing multi-level nuances for real-world challenges in table evaluation resembling humans.
4.2 Human Ranking Correlation Study
To further validate the robustness of our evaluation framework, we conducted a human ranking correlation study using outputs from TabXBench. In this study, expert annotators ranked the quality of a ground-truth table and its perturbed variants each reflecting real-world errors such as structural, semantic, and formatting issues. These human rankings serve as our gold standard for assessing table quality.
For each table, we computed an aggregated score (based on cell-level f1 measures) using various evaluation metrics, including deterministic baselines (e.g., Exact Match, chrf, ROUGE-L) and non-deterministic methods (e.g., BERTScore, H-Score, P-Score, BLUERT, TabEval), alongside our proposed TabXEval. Further, for baselines implementation we run both a single-step and multi-step LLM baseline to populate our proposed rubric tables that mirrors our two-stage TabXEval pipeline more closely along with LLM based ranking and multi-step LLM baseline with Chain-of-thoughts. We then measured the correlation between the automatic rankings and the human judgments using multiple metrics: Spearman’s $\rho$, Kendall’s $\tau$, Weighted Kendall’s $\tau^{\dagger}$, Rank-Biased Overlap (RBO), and Spearman’s Footrule. These measures collectively assess both the overall ranking order and positional differences.
Metrics Spearman’s Kendall’s W-Kendall’s RBO Spearman’s
$\rho$uparrow $\tau$uparrow $\tau^{\dagger}$uparrow $\uparrow$ Footrule $\downarrow$
EM 0.18 0.16 0.16 0.26 0.57
chrF 0.12 0.11 0.08 0.25 0.59
H-Score 0.14 0.11 0.09 0.28 0.51
BERTScore 0.19 0.15 0.13 0.25 0.57
ROUGE-L 0.21 0.18 0.40 0.27 0.53
BLEURT 0.29 0.25 0.25 0.27 0.51
TabEval -0.04 -0.04 -0.03 0.23 0.63
P-Score 0.30 0.27 0.24 0.31 0.39
LLM rubric 0.23 0.16 0.17 0.28 0.47
LLM ranking 0.29 0.24 0.23 0.30 0.41
Multi-prompt 0.29 0.24 0.23 0.30 0.42
Multi-prompt + CoT 0.30 0.25 0.24 0.29 0.45
TabXEval 0.44 0.40 0.38 0.34 0.29
Table 2: Correlation between automatic rankings and human judgments. Higher Spearman’s $\rho$, Kendall’s $\tau$, Weighted Kendall’s $\tau^{\dagger}$, and RBO values indicate better agreement, while lower Spearman’s Footrule values are preferable.
As shown in Table 2, TabXEval achieves the strongest correlation with human rankings across all metrics. Specifically,
Overall Ranking Order: TabXEval attains a Spearman’s $\rho$ of 0.44 a relative improvement of nearly 47% over the next-best method (P-Score at 0.30). Its Kendall’s $\tau$ of 0.40 and Weighted Kendall’s $\tau^{\dagger}$ of 0.38 further indicate strong monotonic agreement with human assessments.
Top-Weighted Agreement: With an RBO of 0.34, TabXEval demonstrates superior alignment in the higher-ranked items, compared to values ranging from 0.23 to 0.31 for other methods.
Positional Accuracy: TabXEval records the lowest Spearman’s Footrule distance (0.29), reflecting minimal positional discrepancy relative to the human gold standard.
Notably, the TabEval method not only fails to capture these nuances as it negatively correlates, highlighting its inability to account for the multifaceted nature of table quality. This analysis reinforces the importance of our two-phase approach, disentangling structural alignment (TabAlign) from detailed cell-level comparison (TabCompare) to effectively mirror human judgment. Finally, the human ranking correlation study clearly demonstrates that TabXEval provides a more robust, interpretable, and human-aligned evaluation of table outputs, capturing both coarse and fine-grained discrepancies that are critical in real-world scenarios.
Figure 3: Sensitivity–Specificity trade-off across metrics. Bubble size reflects the harmonic mean of accuracy, sensitivity, and specificity; color shows F1-score. The dashed line marks ideal balance—top-right methods perform best. Green denotes optimal (Goldilocks) zone; pink favors sensitivity, yellow favors specificity.
4.3 What Sets TabXEval Apart?
A key criteria of any evaluation metric is to achieve a balance between specificity (i.e., avoiding false positives) and sensitivity (i.e., avoiding false negatives). In Figure 3, we visualize this trade-off by plotting each metric’s specificity (y-axis) against its sensitivity (x-axis). The background colormap in the figure corresponds to the F1 score. Finally, Bubble Size represented by harmonic mean of specificity, sensitivity and accuracy, providing a quick visual cue for overall performance.
Goldilocks Zone for Ideal Metrics. Metrics positioned in the top-right portion of the chart (the green-shaded “Goldilocks zone”) demonstrate the desired trait of consistently identifying correct table content (high sensitivity) while minimizing the likelihood of falsely flagging errors (high specificity). TabXEval resides firmly in this zone, illustrating its balanced performance across diverse table perturbations.
Comparisons with Other Metrics.
  • P-Score sits near the Goldilocks zone, indicating solid table-level correctness. However, it lacks explainability and granular insights, offering only a single 0010101010 score limiting interpretability and error traceability.
  • H-Score and BERTScore better capture semantics than string-based metrics (EM, ROUGE-L, chrF), but often miss structural errors like swapped columns or missing rows leading to moderate sensitivity but poor specificity.
  • TabEval relies on entailment over LLM-generated atomic statements, often missing fine-grained numeric/unit mismatches. It showed false positives on re-formatted but semantically correct tables in our tests.
Why TabXBench Matters. TabXBench introduces diverse table perturbations (e.g., missing rows/columns, numeric/unit mismatches) to stress-test metrics across real-world errors (Figure 2). Simple metrics like Exact Match fail under reordering or semantic shifts, while LLM/embedding-based methods miss unit mismatches or partial errors. TabXEval’s two-phase approach first aligning structure (TabAlign), then systematically comparing content (TabCompare) ensures precise discrepancy detection aligned with human judgment.
Significance of TabXEval. By balancing sensitivity, specificity, and F1 scores, TabXEval not only outperforms across evaluation dimensions but also explains its judgments via structured rubrics. This transparency is crucial for financial reporting, scientific validation, and knowledge curation, where subtle errors can be costly. Identifying what went wrong and where, TabXEval provides both quantitative and qualitative insights for improving table generation.
In summary, TabXEval’s interpretability, robust performance, and alignment with TabXBench’s challenging setup establish it as the new standard for explainable, human-aligned table evaluation.
LLaMA-3.3 70B GPT-4o Gemini-2.0-flash
Stat Num String Bool Date List Time Others Num String Bool Date List Time Others Num String Bool Date List Time Others
WikiTables
EI 0.05 4.33 0.00 0.17 0.00 0.00 0.13 0.03 1.17 0.00 0.02 0.00 0.00 0.13 0.02 2.33 0.00 0.11 0.00 0.00 0.07
MI 0.01 0.80 0.00 0.03 0.00 0.00 0.00 0.01 0.93 0.00 0.00 0.00 0.00 0.01 0.00 0.73 0.00 0.00 0.00 0.00 0.00
Partial 0.22 25.00 0.00 0.35 0.00 0.01 0.09 0.32 20.34 0.00 0.55 0.00 0.02 0.10 0.30 22.50 0.00 0.48 0.00 0.02 0.07
WikiBio
EI 0.04 2.84 0.00 0.09 0.00 0.00 0.12 0.04 2.02 0.00 0.03 0.00 0.00 0.09 0.04 2.30 0.00 0.07 0.00 0.00 0.06
MI 0.02 0.29 0.00 0.00 0.00 0.00 0.00 0.00 0.38 0.00 0.01 0.00 0.00 0.00 0.00 0.37 0.00 0.01 0.00 0.00 0.00
Partial 0.16 14.38 0.00 2.60 0.00 0.00 0.03 0.16 15.80 0.00 0.86 0.00 0.00 0.03 0.15 13.59 0.00 2.97 0.00 0.00 0.04
TANQ
EI 0.05 0.84 0.00 0.16 0.09 0.00 0.00 0.02 0.18 0.00 0.06 0.03 0.01 0.00 0.00 0.29 0.00 0.07 0.02 0.00 0.00
MI 0.01 0.24 0.00 0.11 0.01 0.04 0.00 0.01 0.08 0.00 0.05 0.02 0.01 0.00 0.02 0.21 0.00 0.05 0.00 0.02 0.00
Partial 2.73 20.50 0.00 4.72 4.82 2.19 0.07 1.28 11.92 0.00 3.48 3.46 1.32 0.01 1.22 9.35 0.00 2.02 2.64 1.12 0.02
RotoWire
EI 0.84 0.50 0.00 0.00 0.00 0.00 0.01 0.68 0.23 0.00 0.00 0.00 0.00 0.00 1.31 0.54 0.00 0.00 0.00 0.00 0.09
MI 0.97 0.32 0.00 0.00 0.00 0.00 0.04 0.56 0.08 0.00 0.00 0.00 0.00 0.00 1.06 0.28 0.00 0.00 0.00 0.00 0.02
Partial 0.66 0.92 0.00 0.00 0.00 0.00 0.00 0.31 0.69 0.00 0.00 0.00 0.00 0.00 2.72 3.87 0.00 0.00 0.00 0.00 0.03
Table 3: Cell-Level Performance Analysis of Extra (EI), Missing (MI), and Partial mismatches across data types — numerical, string, boolean, date, list, time, and other — for WikiTables, WikiBio, TANQ, and RotoWire. Highlights: GPT-4o shows fewer string EI in WikiTables and lower partial errors in numerical and string cells in TANQ and RotoWire.
4.4 Performance Analysis
The results from our Endurance Test on Text-to-Table Generation as depicted in Table 4 and Table 3 clearly demonstrate how TabXEval’s two-phase evaluation framework enables us to drill down from overall table structure to fine-grained cell details.
Table-Level Performance Across datasets, GPT-4o frequently exhibits higher exact match scores (EM) for both rows and columns compared to LLaMA-3.3 and Gemini-2.0-flash. For instance, on WikiTables, GPT-4o achieves a row EM score of 27.11, surpassing the performance of the other models. Additionally, GPT-4o consistently maintains low Extra Information (EI) values at the column level (e.g., only 0.07 EI on WikiTables), indicating that it preserves the intended table structure with minimal unintended additions. Such metrics underscore the model’s ability to capture overall table integrity across diverse datasets, including WikiBio, TANQ, and RotoWire.
Cell-Level Performance A closer examination at the cell level reveals further nuances. In datasets such as WikiTables and WikiBio, GPT-4o records significantly fewer errors in string cells; for example, its string EI on WikiTables is only 1.17 compared to 4.33 for LLaMA-3.3. On TANQ, both GPT-4o and Gemini-2.0-flash show lower partial errors in numerical and string cell types relative to LLaMA-3.3, suggesting more robust semantic and syntactic matching. Notably, on RotoWire, GPT-4o also demonstrates lower partial error counts in both numerical and string cells when compared with Gemini-2.0-flash. These detailed cell-level insights are crucial as they highlight the models’ abilities to handle fine-grained discrepancies such as unit mismatches or subtle formatting errors. This course-to-fine grain evaluation not only facilitates the identification of specific error types but also offers interpretable insights into the strengths and weaknesses of each model.
Figure 4: Human Ranking Correlation. This plot compares existing metrics (with the best performance indicated by the red dashed line) against various configurations of our TabXEval metric. The green star highlights the best performing TabXEval configuration.
Adaptability and Robustness Evaluation Figure 4 illustrates that TabXEval consistently outperforms existing metrics in human-correlation across a broad range of weighting schemes. Each box plot corresponds to 26 permutations of weights for key dimensions (e.g., missing, extra, row, column, cell, partial) these weights contribute to the Equations 3.2 &  3.2, set to either perm(0)=$\{0,1\}$ or perm=(0.25)$\{0.25,1\}$. This flexibility lets TabXEval adapt to dataset-specific priorities (e.g., penalizing missing rows more heavily) or maintain robust, domain-agnostic performance. Notably, our best-performing configuration (marked by the star) demonstrates both high average correlation and low variance, underscoring TabXEval’s capacity to balance fine-grained descriptors with overall structural fidelity.
5 Conclusion
In this work, we have introduced TabXEval an eXhaustive and eXplainable, two-phase framework that transforms table evaluation by disentangling structural alignment from detailed cell-level comparison. Our method leverages a comprehensive rubric to quantify both coarse and fine-grained errors, yielding results that strongly correlate with human assessments. By developing TabXBench, a challenging multi-domain benchmark with diverse perturbations, we have demonstrated the robustness, explainability, and human-alignment of our approach. While limitations such as computational overhead and handling of hierarchical tables remain, the promising performance of TabXEval opens avenues for further research and refinement in automatic table evaluation.
6 Limitations
While TabXEval demonstrates strong performance, it is not without its limitations. Its reliance on large-scale language models comes at the cost of increased computational overhead, which may impact scalability. Moreover, the method faces challenges when dealing with hierarchical tables, where nested headers and multi-level groupings make alignment significantly more complex. Lastly, as a reference-based evaluation approach, TabXEval necessitates access to ground-truth tables, leaving the question of referenceless evaluation an open and compelling challenge for future research.
7 Ethics Statement
The authors affirm that this work adheres to the highest ethical standards in research and publication. Ethical considerations have been meticulously addressed to ensure responsible conduct and the fair application of computational linguistics methodologies. Our findings are aligned with experimental data, and while some degree of stochasticity is inherent in black-box Large Language Models (LLMs), we mitigate this variability by maintaining fixed parameters such as temperature, $top_{p}$, and $top_{k}$. Furthermore, our use of LLMs, including GPT-4o, Gemini, and Llama, complies with their respective usage policies. To refine the clarity and grammatical accuracy of the text, AI based tools such as Grammarly and ChatGPT were employed. Additionally, human annotators who are also among the authors actively contributed to data labeling and verification, ensuring high-quality annotations. To the best of our knowledge, this study introduces no additional ethical risks.
Acknowledgements
We thank the Complex Data Analysis and Reasoning Lab at Arizona State University for computational support, and Krishna Singh Rajput for his support during the submission process.
Appendix
Appendix A Performance Analysis
LLaMA-3.3 70B GPT-4o Gemini-2.0-flash
MI EI EM MI EI EM MI EI EM
WikiTable
Row 8.69 15.82 25.59 23.44 11.51 27.11 20.81 10.67 26.03
Col 0.37 0.92 1.67 4.47 0.07 1.02 2.97 0.37 1.55
WikiBio
Row 25.16 29.33 16.17 26.09 27.63 19.39 30.08 24.38 16.89
Col 0.10 0.0 0.05 0.05 0.025 0.0 0.12 0.0 0.0
TANQ
Row 7.6 5.83 10.97 8.27 2.80 13.00 8.51 4.01 13.01
Col 2.69 1.82 23.89 2.24 0.19 22.42 2.82 0.63 21.78
RotoWire
Row 3.48 39.22 17.62 1.32 28.65 38.41 3.10 22.82 13.80
Col 10.87 16.21 52.70 17.57 5.71 60.24 16.35 10.52 48.51
Table 4: Table-Level Performance Analysis: Row/Column MI, EI, and EM rates on WikiTables, WikiBio, TANQ, and RotoWire. Highlights: GPT-4o leads with highest Row EM (27.11) and lowest Col EI (0.07) on WikiTables.
Method Spearman’s $\rho$ Kendall’s $\tau$ Weighted Kendall’s $\tau$ RBO Spearman’s Footrule
P-Score (GPT-4o) 0.30 0.27 0.24 0.31 0.39
BleuRT 0.29 0.25 0.25 0.27 0.51
TabXEval (GPT-4o) 0.44 0.40 0.38 0.34 0.29
TabXEval (LLaMA3.3 70B-Instruct) 0.37 0.30 0.30 0.29 0.44
TabXEval (Qwen72B-Instruct) 0.33 0.27 0.28 0.30 0.38
TabXEval (Gemini 2.0 pro) 0.30 0.23 0.21 0.30 0.40
Table 5: Robustness of TabXEval across LLM back-bones. Human-ranking correlations for TabXEval run with GPT-4o, LLaMA-3.3-Instruct, Qwen-72B-Instruct and Gemini-2.0-pro, compared to BLEURT and the P-Score baseline. TabXEval + GPT-4o attains the strongest alignment (Spearman’s $\rho$ = 0.44), but all back-bones retain a $\geq$ 0.30 correlation, indicating backbone-agnostic reliability.
Figure 5: Human Ranking Correlation, existing metric vs ours all configurations.
Appendix B Examples
Below is a detailed breakdown of our scoring method. For example we are evaluating two movie examples where the Ground Truth Table (GT): 5x5 table and a Reference Table (Ref): 4x6 table.
$\mathrm{\textbf{{TabXEval}{}}}=\sum_{I\,\in\,\{\mathrm{Missing},\,\mathrm{Extra},\,\mathrm{Partial}\}}\beta_{I}\Biggl{(}\sum_{E\,\in\,\{\mathrm{row},\,\mathrm{column},\,\mathrm{cell}\}}\alpha_{E}\,\frac{f_{E}}{N_{E}}\Biggr{)}\times\gamma_{p}$ (1)
Partial Score Weight Definition
$\gamma_{p}=\begin{cases}1,&\text{if no partial cell,}\\[6.0pt] \omega_{p}\times\displaystyle\left|\frac{GT-Ref}{Ref}\right|,&\text{if partialdetected.}\end{cases}$
Observed Errors
  • Missing Row: The reference table is missing one row compared to the ground truth.
  • Extra Column: The reference table has an additional column about directors.
  • Partial Match in a Cell: One cell in the reference table (release date) has a typo and increases the date by 2 days.
Weights
Entity Weights
  • Column: $\alpha_{\mathrm{column}}=1$
  • Row: $\alpha_{\mathrm{row}}=0.9$
  • Cell: $\alpha_{\mathrm{cell}}=0.8$
Information Error Weights
  • Missing: $\beta_{\mathrm{Missing}}=1$
  • Extra: $\beta_{\mathrm{Extra}}=0.9$
  • Partial: $\beta_{\mathrm{Partial}}=0.8$
Gamma Modifier for Partial Matches
For numerical data, we use a weight $\omega_{p}=0.9$ and compute a normalized difference:
[…]
Thus,
$\gamma_{p}=0.9\times 0.4=0.36.$
Our proposed rubric is domain agnostic, the beauty of our work is that our weighting scheme is flexible which makes the evaluation metric domain specific based on domain knowledge, for instance one can set higher weights to cell values with numeric types for financial data, in contrast the same rubric can be tuned for sports data where structural nuances should be penalized. Moreover, evidenced by our findings (Figure 4), highlights the strength and generalizability of the underlying scoring mechanism.
Scoring:
Missing Row:
  • Weight for rows: $\alpha_{E}=0.9$
  • Total rows in GT: 5
  • Weight for missing: $\beta_{I}=1$
Contribution to $\text{Contribution to error}=\alpha_{E}\times\beta_{I}\times\frac{\text{Numberof missing rows}}{\text{Total rows}}=0.9\times 1\times\frac{1}{5}=0.18$
Extra Column:
  • Weight for columns: $\alpha_{E}=1$
  • Weight for extra: $\beta_{I}=0.9$
  • Total columns in GT: 5
Contribution to $\text{Contribution to error}=\alpha_{E}\times\beta_{I}\times\frac{1}{5}=1\times 0.9\times\frac{1}{5}=0.18$
Partial Match (Cell):
  • Weight for cells: $\alpha_{E}=0.8$
  • Weight for partial: $\beta_{I}=0.8$
  • Total cells in GT: $5\times 5=25$
  • $\gamma_{p}=0.36$
Contribution to $\text{Contribution to error}=\alpha_{E}\times\beta_{I}\times\frac{1}{25}\times\gamma_{p}=0.8\times 0.8\times\frac{1}{25}\times 0.36=0.0088$
Total Error Score:
$0.18+0.18+0.0088=\boxed{0.368}$
By adjusting the weights assigned to entity types (rows, columns, cells), information errors (missing, extra, partial), and partial matches, the framework can be fine-tuned to emphasize aspects critical to particular applications.
Appendix C Prompts
We provide detailed prompts for TabAlign, TabCompare and Direct-LLM Baseline.
Figure 6: Prompt for tabular alignment, leveraging Partially Aligned Table and Reference Tables to generate a final structured table while preserving unmatched elements.
Figure 7: Prompt for identifying data type, entity, and unit differences between two tables, outputting structured tuples to capture variations in numerical, string, date, and categorical values.
Figure 8: Baseline comparison prompt for evaluating differences between ground truth (GT) and generated data, providing structured metrics for row, column, and cell-level analysis.
Figure 9: Sample perturbed tables from the TabXBench benchmark, illustrating domain-specific corruptions at different difficulty levels. (a) Movie domain with Easy: “Easy” perturbations applied to a clean movie-metadata table, including minor spelling errors in film titles, superficial header rephrasing, simple date-format conversions (e.g., “March 3, 2020” $\leftrightarrow$ “03/03/2020”), trivial numeric formatting changes (addition/removal of thousands separators), and basic unit shifts (e.g., runtime in minutes vs. hours). (b) Finance domain with (Easy + Hard): A financial report table subjected to both “Easy” (currency-symbol normalization, decimal rounding) and “Hard” modifications, such as inconsistent metric abbreviations (e.g., “Rev.” vs. “Revenue”), merged indicator columns, omitted quarterly rows, and large-scale unit mismatches (millions vs. billions). (c) Sports domain with Medium : A sports-stats table with “Medium” perturbations, featuring moderate header reordering (e.g., swapping “Team” and “Position”), slight numeric shifts in game statistics (win/loss counts adjusted by one or two), merged athlete performance rows, and partial row/column transpositions to emulate realistic table-generation errors.